OpenStudy (amistre64):

A boat starts at the equator and travels 1035.835 miles due east and turns north, it then travels 690.557 miles. How far is it away from where it started? Prove your case :)

OpenStudy (kinggeorge):

What number are we using for circumference of the earth?

OpenStudy (amistre64):

i tried to conform it to a radius of earth to be about 3956.6 miles

OpenStudy (kinggeorge):

In that case, the boat traveled about 1/24 of the way around the world to the east, and 1/36 of the way around to the north.

OpenStudy (amistre64):

so far so good, 15 degrees east and 10 degrees north

OpenStudy (amistre64):

i tried for 10 degree; you sure that 1/36 ?

OpenStudy (amistre64):

lol, 10/360 got it lol

OpenStudy (amistre64):

so if we can determine the angle that it is from the start we can find a distance in my view

OpenStudy (kinggeorge):

Just reminding myself of how to find lengths of chords.

OpenStudy (amistre64):

could always try law of cosines to find a missing length on iso tris if your going for a linear measure

OpenStudy (amistre64):

e^2 = r^2+r^2-2r^2 cos(15) = cos(15)

OpenStudy (amistre64):

e = sqrt(cos(15)); n=sqrt(cos(10)) would the chords to me

OpenStudy (kinggeorge):

So after calculating the chords, and then using the pythagorean theorem, I'm getting about 1242 miles.

OpenStudy (amistre64):

my chords are off, so i hope you dint use them :) e = sqrt(2r^2(1-cos(E)))

OpenStudy (kinggeorge):

I was using \[l=2R \sin(\theta/2)\]

OpenStudy (kinggeorge):

where \(l\) is the chord length, \(R\) is the radius, and \(\theta\) is the angle in radians.

OpenStudy (amistre64):

e^2 + n^2 = d^2 2r^2(1-cos(15)) + 2r^2(1-cos(10)) = d^2 sqrt(2r^2(1-cos(15)) + 2r^2(1-cos(10))) = d; 1241 or so is a good measure of linear distance

OpenStudy (amistre64):

how far is it across the surface :)

OpenStudy (kinggeorge):

Using 1242 as the direct distance, we need to reverse the chord length formula to find the angle. So we get \[1242=2*3956.6\sin(\theta/2)\]This implies that \(\theta \approx .1576\) in radians once again. Hence, the arc length is \[L=0.1576(3956.6)\approx623.5\]

OpenStudy (amistre64):

ill accept that ;) good job

OpenStudy (amistre64):

it should be greater than 1035 miles tho right?

OpenStudy (kinggeorge):

Thank you. I was rounding to 3 or 4 decimal places fairly often, so that likely introduced some error.

OpenStudy (amistre64):

i think my way get me about 1247 miles away on the surface

OpenStudy (kinggeorge):

Hmmm. That is a good point. I would expect it to be larger than 1035. Let me see if I can find where I went wrong.

OpenStudy (kinggeorge):

Herp derp. I was solving for \(\theta/2\) not \(\theta\). So multiply my answer by 2 to get \(2\cdot623.5=1247\) as the actual arc length.

OpenStudy (amistre64):

:) thats better